OpenStudy (anonymous):

x+3y=9 graph using slope and y-intercept

OpenStudy (anonymous):

Are you literally copy+pasting questions from your test?

Parth (parthkohli):

http://wolframalpha.com or your human brain. Choose amongst them ;)

OpenStudy (anonymous):

at vf321, no

OpenStudy (anonymous):

@ParthKohli, thanks! I am always interested in websites that will help

Parth (parthkohli):

You're welcome, but I really prefer the latter for practice.

OpenStudy (anonymous):

i agree

Parth (parthkohli):

This is a linear equation--which makes a straight line. What you have to do is fetch two points from which the line will pass through(or convert this equation into the standard equation of a straight line).

OpenStudy (anonymous):

\[x+3y=9\] \[3y=-x+9\] \[y=-\frac{1}{3}x+3\] but it would be easier to graph \[x+3y=9\] by finding the \(x\) and \(y\) intercepts, plotting the two points and connecting the dots

Parth (parthkohli):

Then some kinda thing you have to remember: 1) y-intercept is y when x = 0. 2) slope is the change in y per change in x.

OpenStudy (anonymous):

@ParthKohli will show you how to find them, it is the easiest way in this example

OpenStudy (anonymous):

that is what I need to fi do is get the points to graph

mathslover (mathslover):

first of all find the x intercept : by putting y = 0 what do you get ?: x+3(0) = 9

Parth (parthkohli):

Okay sir satellite. 1) Find \(x\) when \(y = 0\). This is your \(x\)-intercept. 2) Find \(y\) when \(x = 0\). This is your \(y\)-intercept.

OpenStudy (anonymous):

x=9

mathslover (mathslover):

good now find y intercept by putting x = 0 0 + 3y =9

Parth (parthkohli):

That's right. The \(x\)-intercept is (9,0).

OpenStudy (anonymous):

y=3

mathslover (mathslover):

solve for y : 0 + 3y = 9 tell me what you get?

Parth (parthkohli):

Righty right!

mathslover (mathslover):

good that is (0,3)

mathslover (mathslover):

now can u plot them on graph?

Parth (parthkohli):

(9,0) and (0,3) are two points that the line will pass through, okay?

mathslover (mathslover):

the graph will be like this

Parth (parthkohli):

Just make these points, and make a straight line through them.

OpenStudy (anonymous):

thank you all for simplifying

Parth (parthkohli):

You're welcome! I always like to help... :)

mathslover (mathslover):

Welcome and best of luck for your journey here and there (slope-intercept)

OpenStudy (anonymous):

can you all help me with another more complicated problem?

mathslover (mathslover):

ok

mathslover (mathslover):

post that as a new question

Parth (parthkohli):

All right. Please try some of it by yourself.

OpenStudy (anonymous):

(-3,-1) slope is 2/3. I need to graph